OpenStudy (anonymous):

sqrt(x+16) - sqrt(x) - 2 = 0

OpenStudy (anonymous):

x=25 is what I got

OpenStudy (jdoe0001):

how did you get 25?

OpenStudy (anonymous):

I figured it out is it right

OpenStudy (jdoe0001):

how did you figure it was right?

OpenStudy (anonymous):

Sub it back in to check.

OpenStudy (anonymous):

I solved it Using order of operations and I subbed it in I just need a second opinion please.

OpenStudy (jdoe0001):

using the order of operations? that's a lnear simplification, not an operational equation

OpenStudy (anonymous):

Ok I really need a straight answer someeone I already figured it out I did the work

OpenStudy (anonymous):

\[\sqrt{41}-\sqrt{25}-2=-0.596...\ne0\]

OpenStudy (anonymous):

Plug in 25 for x √(25)-16 = √(25)-2

rebeccaxhawaii (rebeccaxhawaii):

x=9

OpenStudy (anonymous):

OK can you show me the right way

OpenStudy (jdoe0001):

it doesn't quite factor nicely.... so... can you post a screenshot of the material?

OpenStudy (jdoe0001):

hmm in short, she didn't do it

rebeccaxhawaii (rebeccaxhawaii):

yeah pretty much but i got the steps to further her understanding

OpenStudy (jdoe0001):

hmmm kinda like a blind leading a blind :/

rebeccaxhawaii (rebeccaxhawaii):

can't argue with a calculator

rebeccaxhawaii (rebeccaxhawaii):

tbh i helped her more than you

OpenStudy (anonymous):

Thanks rebecca and jdoe please stop because you did nothing but be obnoxious

OpenStudy (jdoe0001):

anyhow 9 doesn't check out I was checking about how to factor.... doesn't quite nicely

rebeccaxhawaii (rebeccaxhawaii):

no problem we jdoe i check through 2 different calcs

rebeccaxhawaii (rebeccaxhawaii):

checked *

rebeccaxhawaii (rebeccaxhawaii):

2 calculators > you

OpenStudy (jdoe0001):

hmmm well... yes... I'm somewhat sure the calculator is ok but that simply means, the credit goes to the calculator, not you either way, you're simply back to square one but I was looking at it... and anyway... so it goes

OpenStudy (anonymous):

I got the answer right... lol jdoe you are making a fool of yourself rn

OpenStudy (anonymous):

LOL he blocked me right now

OpenStudy (anonymous):

\[\sqrt{x+16}=\sqrt{x}+2\] squaring both sides \[x+16=x+4+4\sqrt{x}\] \[4\sqrt{x}=16-4=12,\sqrt{x}=3,x=9\]